at what point do the graphs of the equations y=7x-3 and -6x+y=2 intersect
Answer by mananth(16946)

equation1
when x= 0,y=-3, (0,-3 )
when y =0,x=0.43,(0.43,0)
Equation 2
when x= 0,y=2(0,2)
when y =0,x=-0.33(-0.33,0)

the lines intersect at (5,32)
7 x -1 y = 3 .............1
-6 x + 1 y = 2 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
7 x 1 y = 3
-6 x 1 y = 2
Add the two equations
1 x = 5
/ 1
x = 5
plug value of x in (1)
7 x + -1 y = 3
35 + -1 y = 3
-1 y = 3 -35
-1 y = -32
y = 32

m.ananth@hotmail.ca